What is the molarity of an HCl solution iſ 28.5 mL of 0.16 M NaOH are needed to neutralize 15.0 mL of the sample? MHCI
Answer #1

The balanced equation is

HCl + NaOH -----> NaCl + H2O

Number of moles of NaOH = molarity * volume of solution in L

Number of moles of NaOH = 0.16 * 0.0285 = 0.00456 mole

From the balanced equation we can say that

1 mole of NaOH requires 1 mole of HCl so

0.00456 mole of NaOH will require

= 0.00456 mole of NaOH *(1 mole of HCl / 1 mole of NaOH)

= 0.00456 mole of HCl

molarity of HCl = number of moles of HCl / volume of solution in L

molarity of HCl = 0.00456 / 0.015 = 0.304 M

Therefore, the molarity of HCl = 0.304 M